OpenStudy (anonymous):

what is the derivative of 3, x^2,X^3,X^5,X, 7x?

OpenStudy (anonymous):

find the derivative by using the formual \[nx ^{n-1}\]

OpenStudy (anonymous):

n is the exponent

OpenStudy (anonymous):

okay, let me find out and i will ask you if its correct or not.

OpenStudy (anonymous):

X^2+= 2x X^3= 3x square X^5= 5x^4 x= 1/x 7x= 7 ??

OpenStudy (baldymcgee6):

The derivative of x is 1

OpenStudy (anonymous):

yes thats right, except for x, it doesnt equal 1/x, it equals 1

OpenStudy (anonymous):

derivative of x is 1, other answers look good

OpenStudy (anonymous):

How do you get that 1?

OpenStudy (anonymous):

because x is actually the same as x^1 and x^(1 - 1) = x^0 and anything raised to the power of 0 is 1

OpenStudy (anonymous):

thats why 7x becomes just 7 because it turns int 7* x^0 which is just 7*1 = 7

OpenStudy (anonymous):

yeah, i got it now.

OpenStudy (anonymous):

Thank you so much.